# Learning to Drive Ethically: Moral Reasoning in Autonomous Driving

### v2

@@ 3.3 Collision probabilities
To estimate the potential risk to other road users, we begin by computing the collision probabilities between the ego vehicle and its surrounding road users. Since the framework optimizes trajectories over varied time horizons, these probabilities are evaluated at each discrete timestep within that horizon. Given the planned trajectory of the ego vehicle, the next step is to predict the future trajectories of nearby road users. For surrounding vehicles, we use an LSTM-based neural network that incorporates convolutional social pooling and semantic road context, using bird’s-eye-view images as input (Geisslinger et al., 2021). It provides at each timestep $\tau$ a mean position $\mu_{o,\tau}$ and a covariance matrix $\mathbf{\Sigma}_{o,\tau}$, defined over the longitudinal distance along the path and the lateral displacement. Under the Gaussian assumption, these parameters provide the probability distribution of the predicted positions of the agent. Additional implementation details are provided in (Geisslinger et al., 2021). For other types of agents, such as pedestrians and cyclists, ground-truth trajectories are used with Gaussian noise added to capture uncertainty. The predicted trajectory of agent $o$ is thus represented as a sequence of Gaussian variables,
@@
Unlike prior work that directly estimates collision probabilities, our method adopts a two-stage approach. In the first stage, we apply the Separating Axis Theorem (SAT) to check for potential overlaps between the ego vehicle and other agents, based on their mean trajectories Gottschalk et al. (1996). This check is performed at every timestep of the planning horizon. If an overlap is detected, we proceed to the second stage, where we compute a more precise proximity-based collision probability using the Mahalanobis distance. This two-stage approach improves efficiency by skipping expensive computations for clearly separated agents.
@@
Under the assumption of Gaussianity, the squared Mahalanobis distance follows a chi-squared distribution with 2 degrees of freedom, i.e., $(D_{m,\tau})^{2}\sim\chi^{2}(2)$. We convert the distance into a collision probability using the survival distribution function of the chi-squared distribution:
@@ 3.4 Risk calculation
In this work, the risk $R_{\tau}$ associated with a traffic participant at time $\tau$ is defined as the expected collision risk with other traffic participants. This risk is calculated as the product of the collision probability $P^{\tau}$} and the estimated harm $H^{\tau}$, i.e., $R_{\tau}=P^{\tau}H^{\tau}$. While the collision probabilities are obtained from Section 3.3, we estimate the harm using empirical equations proposed in (Geisslinger et al., 2023). The harm $H$ between two traffic participants, $A$ and $B$, is modeled as:
@@ 3.6 Prioritized experience replay
with the error ratio $r_{e}=\delta_{r,i}/\delta_{c,i}$, clipped to the range [0.2, 5] to ensure training stability. This adaptive prioritization mechanism enables the agent to focus learning on whichever signal, reward or cost, is more informative at a given time. As a result, the policy evolves with a balanced emphasis on both task performance and ethical responsibility. To prove the effectiveness of this dynamic PER, we also perform an ablation study, detailed in Section 5.
@@ 4.1 Simulation Environment
In this work, we use the MetaDrive simulator (Li et al., 2022) as the training and evaluation environment due to its ability to integrate real-world driving data, enabling the simulation of realistic driving scenarios. This feature is crucial for our study, as we aim to assess the ethical decision-making capabilities of agents in settings that closely mirror real-world conditions, thereby enhancing the relevance and reliability of our evaluation. For the real-world data, we utilize the Waymo Open Dataset (Ettinger et al., 2021), which contains recordings from six cities across the United States. It includes a wide variety of challenging scenarios, such as unprotected turns, merges, lane changes, and complex intersection interactions, making it particularly well-suited for our task.
@@ 4.3.1 Path Planner
Following the approach in (Werling et al., 2010), and under the assumptions $\ddot{l}_{T}=0$, $\dot{d}_{T}=0$, and $\ddot{d}_{T}=0$, the resulting trajectory minimizes jerk, thereby enhancing passenger comfort. In this framework, vehicle trajectories are generated to achieve a target longitudinal speed $\dot{l}_{T}$ and a desired lateral position $d_{T}$. To enable the RL agent to adaptively control driving behavior in response to dynamic traffic environments, the planning horizon $T$ is treated as an action output of the agent. This design allows the agent to flexibly determine the duration of each trajectory segment. With the solutions appended in Section A.2 for equation (20), a sequence of waypoints is produced, which are then tracked by a path-following controller. To discourage the generation of uncomfortable or impractical trajectories, a penalty term on longitudinal jerk is incorporated into the reward function, as detailed in Section 4.2. This encourages the agent to generate smooth and context-appropriate trajectories.
@@ 4.3.2 Path Follower
To ensure accurate trajectory tracking, the vehicle is equipped with two controllers: a PID controller for longitudinal control and a Stanley controller (Thrun et al., 2006) for lateral control. The longitudinal controller regulates the speed by continuously minimizing the error between the current and target longitudinal velocities. This error, denoted by $e(\tau)$, is used to compute the required acceleration through a combination of proportional ($P$), integral ($I$), and derivative ($D$) terms. The resulting control law is expressed as:
@@
For lateral control, the steering angle is used as the control output, and a Stanley controller is employed for this purpose. The Stanley method is a geometric path-tracking controller that gained prominence after contributing to the victory of Stanford in the 2006 DARPA Grand Challenge (Thrun et al., 2006). It is particularly effective because it accounts for both lateral displacement and heading errors. The steering command $\delta(\tau)$ generated by the controller is defined as:
@@ 4.4 Baselines
Figure 2: Rewards and costs obtained by different agents during the training process, with the cost mode set to ethical and cost limit set to 1.
Each agent was trained for 1.5 million timesteps across multiple runs using different random seeds. The simulation environment operates with a timestep interval of 0.1 seconds with 1300 real-world scenarios from the Waymo dataset. To evaluate performance under varying safety constraints, we define three different cost limits: $\eta=\{0.6,0.75,1.0\}$. Figure 2 presents the results for all agents trained under the ethical mode with the cost limit set to $\eta=1.0$, showing both the cumulative reward and corresponding cost over time.
From the average reward plots in Figure 2, all agents appear to converge to a similar reward level. However, the average cost plots reveal that only the EthicAR agent successfully converges while satisfying the cost constraint. This indicates that, unlike EthicAR, the baseline agents exhibit limitations in enforcing cost-aware behavior during training. In particular, for EthicAR w/o PER, as the costs are near zero for most of the training steps, the agent has a hard time to learn effective cost-constrained behavior in the absence of dynamic PER. The SACLAG agent performs better in this regard, maintaining a relatively low cost, yet still fails to meet the cost limit consistently. These results highlight the importance of both the LSTM architecture and dynamic PER in enabling the agent to learn under explicit cost constraints.
@@ 5.1 Risk evaluation
First, to evaluate the effectiveness of the ethical considerations in EthicAR and other agents, we logged the risks discussed in Section 3.4 during evaluation. These include the risk to the ego vehicle, referred to as ego risk, and the risk to other traffic participants, referred to as other risk. The results are shown in Figure 3 and Table 2 and 3, where each agent is evaluated in two modes across three different cost limits. An exception is the LSTMSAC agent, which does not require a cost limit for cost reduction during training, therefore, its results remain the same across all three cost limits.
@@
Apart from the risk analysis, we also examine the dynamics of the autonomous vehicle (AV) to evaluate the comfort level of the proposed framework. Specifically, we analyze the accelerations and jerks of EthicAR operating in both ethical and selfish modes, as well as other agents. The results are shown in Figure 4. For acceleration, values below $1\mathrm{m/s^{2}}$ are generally considered comfortable, while values up to $2\mathrm{m/s^{2}}$ are tolerable in more dynamic scenarios. In terms of jerk, comfort thresholds are typically below $1\mathrm{m/s^{3}}$, with values between $1$ and $3\mathrm{m/s^{3}}$ often associated with reduced comfort or instability (De Winkel et al., 2023).
@@ 5.3 Scenario analyses
To provide a detailed comparison of driving behaviors, we selected four real-world scenarios that each pose an ethical dilemma in everyday traffic. We evaluated EthicAR alongside three baseline agents under a uniform cost limit of 0.6. Each simulation advanced in 0.1s timesteps, and for every agent, we identified five critical instants at which we recorded the TTC with the most relevant road user as well as the peak other risk experienced by the ego vehicle. We then plotted TTC and risk side by side to enable clear, quantitative comparisons. In these plots, any TTC below 0s, indicating the two objects are diverging, or above 10s, indicating negligible imminent danger, is represented as inf, marking the absence of meaningful collision risk at that moment. In the following sections, we show two scenarios in turn, highlighting how EthicAR and the baselines navigate these ethical trade-offs. Two more scenarios are appended in Section A.3.
@@
In the first scenario, our ego vehicle follows a cyclist on a two-lane, bidirectional road while encountering dense traffic congestion in the opposing lane. Figure 6 depicts these results: (a) shows real-time TTC to the cyclist of EthicAR and the baseline agents alongside their peak other risk at each timestep, and (b) presents the corresponding TTC and maximum risk values at the five selected timesteps.
The standard-mode agent, trained solely on efficiency and safety objectives with the reward function (17), exhibits its limitations in this scenario. Starting from a TTC to the cyclist of 3.2s and a low-risk situation, the standard mode, like other baselines, adjusts its heading to achieve a TTC of inf by laterally diverging from the cyclist. However, with no ethical constraints, it immediately attempts to overtake once the TTC reaches inf, closing to an uncomfortably small lateral gap. In reality, forcing an overtake on a narrow, two-lane road with oncoming congestion violates both ethical norms and common etiquette (Luetge, 2017), and risks a collision either with the cyclist or opposing traffic due to the unpredictability of the cyclist. Such unsafe aggressiveness cannot be mitigated by other baselines: both the PER-ablated EthicAR and the SACLAG agent display similarly risky overtaking attempts here. By contrast, EthicAR maintains a respectful distance from the cyclist throughout and refrains from overtaking, keeping its maximum other risk at a low 0.1 until the conclusion of the scenario, demonstrating more socially acceptable, ethically aligned behavior suitable for everyday driving. Importantly, despite the cost function not explicitly prioritizing the minimization of risks to VRUs, EthicAR nonetheless exhibited responsible decision-making toward them. This outcome underscores the effectiveness of the proposed framework and further suggests that the protection of VRUs could be enhanced through straightforward adjustments to weight assignments.
Figure 6: Real-world driving scenario in which the ego vehicle (green) follows a cyclist (red) along a two-way road with oncoming traffic (gray). (a) Per-timestep behavior of EthicAR and baseline agents, showing real-time TTC with the cyclist and the maximum other risk of the ego vehicle. (b) TTC versus maximum risk at the five selected instants, plotted for direct quantitative comparison.
@@
In everyday driving etiquette at an unsignalized intersection, especially when the ego vehicle approaches from the right, it technically has priority to proceed first. However, safety considerations lead all agents to active yield. While this reflects legal norms, not all agents minimize risk effectively. For instance, for the standard mode, starting from inf TTC, and risk of 0, and then allows the other vehicle to turn. Yet its approach remains aggressive: by $4\mathrm{s}, TT$C falls to just 0.4s and risk spikes to 0.7, even though no collision occurs. Such behavior is risky and unsuitable for AV deployment. EthicAR w/o PER agent displays a similar pattern. The SACLAG agent demonstrates improved restraint: though ego speed remains relatively high at $4\mathrm{s}, TT$C and risk increase to about 0.3, still yielding more courteously. By contrast, EthicAR maintains the lowest risk level throughout the scenario, embodying responsible and socially considerate driving.
Figure 7: Real-world driving scenario in which the ego vehicle (green) and an opposing left-turning vehicle (red) both prepare to turn across each other, with other traffic shown in gray. (a) Per-timestep behavior of EthicAR and baseline agents, showing real-time TTC with the cyclist and the maximum other risk of the ego vehicle. (b) TTC versus maximum risk at the five selected instants, plotted for direct quantitative comparison.

### v3

@@ 3.3 Collision probabilities
To estimate the potential risk to other road users, we begin by computing the collision probabilities between the ego vehicle and its surrounding road users. Since the framework optimizes trajectories over varied time horizons, these probabilities are evaluated at each discrete timestep within that horizon. Given the planned trajectory of the ego vehicle, the next step is to predict the future trajectories of nearby road users. For surrounding vehicles, we use an LSTM-based neural network that incorporates convolutional social pooling and semantic road context, using bird’s-eye-view images as input [geisslinger2021watch]. It provides at each timestep $\tau$ a mean position $\mu_{o,\tau}$ and a covariance matrix $\mathbf{\Sigma}_{o,\tau}$, defined over the longitudinal distance along the path and the lateral displacement. Under the Gaussian assumption, these parameters provide the probability distribution of the predicted positions of the agent. Additional implementation details are provided in [geisslinger2021watch]. For other types of agents, such as pedestrians and cyclists, ground-truth trajectories are used with Gaussian noise added to capture uncertainty. The predicted trajectory of agent $o$ is thus represented as a sequence of Gaussian variables,
@@
Unlike prior work that directly estimates collision probabilities, our method adopts a two-stage approach. In the first stage, we apply the Separating Axis Theorem (SAT) to check for potential overlaps between the ego vehicle and other agents, based on their mean trajectories [gottschalk1996obbtree]. This check is performed at every timestep of the planning horizon. If an overlap is detected, we proceed to the second stage, where we compute a more precise proximity-based collision probability using the Mahalanobis distance. This two-stage approach improves efficiency by skipping expensive computations for clearly separated agents.
@@
Assuming local Gaussianity [geisslinger2021watch], the squared Mahalanobis distance follows a chi-squared distribution with 2 degrees of freedom, i.e., $(D_{m,\tau})^{2}\sim\chi^{2}(2)$. We convert the distance into a collision probability using the survival distribution function of the chi-squared distribution:
@@ 3.4 Risk calculation
In this work, the risk $R_{\tau}$ associated with a traffic participant at time $\tau$ is defined as the expected collision risk with other traffic participants. This risk is calculated as the product of the collision probability $P^{\tau}$} and the estimated harm $H^{\tau}$, i.e., $R_{\tau}=P^{\tau}H^{\tau}$. While the collision probabilities are obtained from Section 3.3, we estimate the harm using empirical equations proposed in [geisslinger2023ethical]. The harm $H$ between two traffic participants, $A$ and $B$, is modeled as:
@@ 3.6 Prioritized experience replay
with the error ratio $r_{e}=\delta_{r,i}/\delta_{c,i}$, clipped to the range [0.2, 5] to ensure training stability. This adaptive prioritization mechanism enables the agent to focus learning on whichever signal, reward or cost, is more informative at a given time. As a result, the policy evolves with a balanced emphasis on both task performance and ethical responsibility. The overall algorithm pseudocode is shown in Algorithm 1. To provide evidence of the effectiveness of this dynamic PER, we also perform an ablation study, detailed in Section 5.
@@ 4.1 Simulation Environment
In this work, we use the MetaDrive simulator [li2022metadrive] as the training and evaluation environment due to its ability to integrate real-world driving data, enabling the simulation of realistic driving scenarios. This feature is crucial for our study, as we aim to assess the ethical decision-making capabilities of agents in settings that closely mirror real-world conditions, thereby enhancing the relevance and reliability of our evaluation. For the real-world data, we utilize the Waymo Open Dataset [ettinger2021large], which contains recordings from six cities across the United States. It includes a wide variety of challenging scenarios, such as unprotected turns, merges, lane changes, and complex intersection interactions, making it particularly well-suited for our task. During training and evaluation, other vehicles (besides the ego vehicle) are reactive, meaning that they primarily follow the trajectories provided in the Waymo Dataset. However, when a potential rear-end collision is detected, their behavior is replaced with the intelligent driver model (IDM) [treiber2000congested] and MOBIL model [kesting2007general].
@@ 4.3.1 Path Planner
Following the approach in [werling2010optimal], and under the assumptions $\ddot{l}_{T}=0$, $\dot{d}_{T}=0$, and $\ddot{d}_{T}=0$, the resulting trajectory minimizes jerk, thereby enhancing passenger comfort. In this framework, vehicle trajectories are generated to achieve a target longitudinal speed $\dot{l}_{T}$ and a desired lateral position $d_{T}$. To enable the RL agent to adaptively control driving behavior in response to dynamic traffic environments, the planning horizon $T$ is treated as an action output of the agent. This design allows the agent to flexibly determine the duration of each trajectory segment. With the solutions appended in Section A.3 for equation (20), a sequence of waypoints is produced, which are then tracked by a path-following controller. To discourage the generation of uncomfortable or impractical trajectories, a penalty term on longitudinal jerk is incorporated into the reward function, as detailed in Section 4.2. This encourages the agent to generate smooth and context-appropriate trajectories.
@@ 4.3.2 Path Follower
To ensure accurate trajectory tracking, the vehicle is equipped with two controllers: a PID controller for longitudinal control and a Stanley controller [thrun2006stanley] for lateral control. The longitudinal controller regulates the speed by continuously minimizing the error between the current and target longitudinal velocities. This error, denoted by $e(\tau)$, is used to compute the required acceleration through a combination of proportional ($P$), integral ($I$), and derivative ($D$) terms. The resulting control law is expressed as:
@@
For lateral control, the steering angle is used as the control output, and a Stanley controller is employed for this purpose. The Stanley method is a geometric path-tracking controller that gained prominence after contributing to the victory of Stanford in the 2006 DARPA Grand Challenge [thrun2006stanley]. It is particularly effective because it accounts for both lateral displacement and heading errors. The steering command $\delta(\tau)$ generated by the controller is defined as:
@@ 4.4 Baselines
Figure 2: Training rewards and costs of different agents, together with the evolution of Lagrange multipliers in safe RL. Results are obtained with the cost mode set to ethical and the cost limit fixed at 1.
Each agent was trained for 1.5 million timesteps across multiple runs using different random seeds. The simulation environment operates with a timestep interval of 0.1 seconds with 1300 real-world scenarios from the Waymo dataset. To evaluate performance under varying safety constraints, we define six different cost limits: $\eta=\{0.1,0.3,0.6,0.75,1,2\}$, which represent varying levels of strictness. Figure 2 presents the results for all agents trained under the ethical mode with the cost limit set to $\eta=1.0$, showing both the cumulative reward and corresponding cost over time.
From the average reward plots in Figure 2, all agents appear to converge to a similar reward level. However, the average cost plots reveal that only the EthicAR agent successfully converges while satisfying the cost constraint. This indicates that, unlike EthicAR, the baseline agents exhibit limitations in enforcing cost-aware behavior during training. In particular, for EthicAR w/o PER, as the costs are near zero for most of the training steps, the agent has a hard time to learn effective cost-constrained behavior in the absence of dynamic PER. The SACLAG agent performs better in this regard, maintaining a relatively low cost, yet still fails to meet the cost limit consistently.
The evolution of Lagrange multipliers in Figure 2 during training exhibits consistent trends across agents. According to Equation (3), the Lagrange multiplier $\lambda$ regulates the trade-off between reward maximization and constraint satisfaction. For the EthicAR agent, $\lambda$ initially rises as the agent explores and frequently violates cost constraints, thereby enforcing constraint adherence. As training progresses, $\lambda$ oscillates to balance rewards and costs before eventually plateauing at levels that maintain constraint satisfaction without excessive penalization. In contrast, the SACLag agent, while able to detect rare cost violations through dynamic PER, fails to converge to stable constraint levels due to the absence of the LSTM architecture. Consequently, it cannot achieve the desired balance between reward optimization and cost constraint enforcement. For EthicAR w/o PER, the scarcity of cost signals prevents the agent from learning meaningful behaviors related to cost limitation. As a result, $\lambda$ remains near zero throughout training, and the agent fails to develop sensitivity to cost constraints. These findings highlight the critical role of both the LSTM architecture and dynamic PER in enabling effective learning under explicit cost constraints.
@@ 5 Evaluation and results
5.1 Constraint satisfaction
First, we evaluate how well different agents satisfy cost constraints across two modes using a compliance metric, reporting both overall and risk-conditioned compliance. Compliance is defined as the percentage of steps in which the cost remains below the threshold $\eta$. Intuitively, higher compliance indicates that the agent more frequently operates within ethical safety limits. Since constraint enforcement is only meaningful under conditions where violations can occur, we measure compliance in two ways: (1) across all steps, which reflects overall stability over episodes, and (2) restricted to risky steps, where the risk value exceeds zero. We compare our EthicAR agent against three baselines, SACLAG, EthicAR w/o PER, and LSTMSAC, across six constraint thresholds in both ethical and selfish modes. Results are shown in Figure 3.
From Figure 3, we identify $\eta\in[0.6,1]$ as the suitable operating range. Thresholds below 0.6 are unrealistically strict, such that almost no agent can comply, while thresholds above 1 are too permissive to meaningfully enforce ethical behavior. In practice, the interval $[0.6,1]$ balances feasibility and enforcement, reflecting scenarios where the agent must reliably avoid severe violations while tolerating small, unavoidable costs due to environmental uncertainty. Accordingly, we focus on cost thresholds $\eta=\{0.6,0.75,1\}$ in the following analysis.
Across both modes, EthicAR consistently achieved the highest compliance, exceeding $88\sim 96\%$ on all steps and $90\sim 95\%$ on risky steps in the suitable threshold range. This indicates that EthicAR not only maintains cost constraints under normal operation but also reliably avoids rare but critical violations. SACLAG achieved moderate compliance ($68\sim 78\%$ on all steps; $60\sim 70\%$ on risky steps), suggesting partial success in constraint learning but limited robustness in rare-event cases. EthicAR w/o PER performed poorly, with compliance below $30\%$ on risky steps, confirming that prioritized experience replay is essential for learning ethical behavior when violations are rare. Finally, the LSTMSAC agent remained at floor levels ($15\%$), highlighting the gap between structured Safe RL and naive behavior. A key observation is that overall compliance can appear inflated due to the predominance of trivially safe steps, whereas risky-step compliance better reflects an agent’s ability to enforce constraints in critical situations. The similarity of results between ethical and selfish modes further shows that these findings are robust across training conditions.
Taken together, these results show that EthicAR successfully enforces ethical constraints even under rare-event conditions, while SACLAG only partially succeeds and alternative approaches fail. The consistency across modes highlights both the generality of these findings and the importance of combining LSTM-based architectures with dynamic PER for safety-critical, ethics-aware learning.
Figure 3: Constraint compliance of Safe RL agents across thresholds and modes. Compliance is shown for four agents (EthicAR, SACLag, EthicAR w/o PER, and LSTMSAC) across six cost thresholds ($\eta=0.1\sim 2$). Results are reported for all steps (right panels) and risky steps only (with a risk greater than 0), under two training modes: ethical (top) and selfish (bottom). EthicAR achieves the highest compliance, particularly in the suitable threshold range ($\eta=0.6\sim 1$), while SACLag learns partially, and EthicAR w/o PER fails to adapt to rare but critical violations. The LSTMSAC agent shows near-zero compliance across all settings. The selfish mode yields similar outcomes, confirming robustness across configurations.
@@ 5.2 Risk evaluation
Then, to further evaluate the effectiveness of the ethical considerations in EthicAR and other agents, we logged the risks discussed in Section 3.4 during evaluation. These include the risk to the ego vehicle, referred to as ego risk, and the risk to other traffic participants, referred to as other risk. The results are shown in Figure 4 and Table 2 and 3, where each agent is evaluated in two modes across three different cost limits. An exception is the LSTMSAC agent, which does not require a cost limit for cost reduction during training, therefore, its results remain the same across all three cost limits.
@@
Apart from the risk analysis, we also examine the dynamics of the AV to evaluate the comfort level of the proposed framework. Specifically, we analyze the accelerations and jerks of EthicAR operating in both ethical and selfish modes, as well as other agents. The results are shown in Figure 5. For acceleration, values below $1\mathrm{m/s^{2}}$ are generally considered comfortable, while values up to $2\mathrm{m/s^{2}}$ are tolerable in more dynamic scenarios. In terms of jerk, comfort thresholds are typically below $1\mathrm{m/s^{3}}$, with values between $1$ and $3\mathrm{m/s^{3}}$ often associated with reduced comfort or instability [de2023standards].
@@ 5.4 Scenario analyses
To provide a detailed comparison of driving behaviors, we selected four real‑world scenarios that each pose an ethical dilemma in everyday traffic. We evaluated EthicAR alongside three baseline agents under a uniform cost limit of 0.6. Each simulation advanced in 0.1s timesteps, and for every agent, we identified five critical instants at which we recorded the TTC with the most relevant road user as well as the peak other risk experienced by the ego vehicle. We then plotted TTC and risk side by side to enable clear, quantitative comparisons. In these plots, any TTC below 0s, indicating the two objects are diverging, or above 10s, indicating negligible imminent danger, is represented as inf, marking the absence of meaningful collision risk at that moment. In the following sections, we show two scenarios in turn, highlighting how EthicAR and the baselines navigate these ethical trade‑offs. Two more scenarios are appended in Section A.4.
@@
In the first scenario, our ego vehicle follows a cyclist on a two‑lane, bidirectional road while encountering dense traffic congestion in the opposing lane. Figure 7 depicts these results: (a) shows real‑time TTC to the cyclist of EthicAR and the baseline agents alongside their peak other risk at each timestep, and (b) presents the corresponding TTC and maximum risk values at the five selected timesteps.
The standard‑mode agent, trained solely on efficiency and safety objectives with the reward function (17), exhibits its limitations in this scenario. Starting from a TTC to the cyclist of 3.2s and a low-risk situation, the standard mode, like other baselines, adjusts its heading to achieve a TTC of inf by laterally diverging from the cyclist. However, with no ethical constraints, it immediately attempts to overtake once the TTC reaches inf, closing to an uncomfortably small lateral gap. In reality, forcing an overtake on a narrow, two‑lane road with oncoming congestion violates both ethical norms and common etiquette [luetge2017german], and risks a collision either with the cyclist or opposing traffic due to the unpredictability of the cyclist. Such unsafe aggressiveness cannot be mitigated by other baselines: both the PER‑ablated EthicAR and the SACLAG agent display similarly risky overtaking attempts here. By contrast, EthicAR maintains a respectful distance from the cyclist throughout and refrains from overtaking, keeping its maximum other risk at a low 0.1 until the conclusion of the scenario, showing more socially acceptable, ethically aligned behavior suitable for everyday driving. Importantly, despite the cost function not explicitly prioritizing the minimization of risks to VRUs, EthicAR nonetheless exhibited responsible decision-making toward them. This outcome underscores the effectiveness of the proposed framework and further suggests that the protection of VRUs could be enhanced through straightforward adjustments to weight assignments.
Figure 7: Real‑world driving scenario in which the ego vehicle (green) follows a cyclist (red) along a two‑way road with oncoming traffic (gray). (a) Per‑timestep behavior of EthicAR and baseline agents, showing real‑time TTC with the cyclist and the maximum other risk of the ego vehicle. (b) TTC versus maximum risk at the five selected instants, plotted for direct quantitative comparison.
@@
In everyday driving etiquette at an unsignalized intersection, especially when the ego vehicle approaches from the right, it technically has priority to proceed first. However, safety considerations lead all agents to active yield. While this reflects legal norms, not all agents minimize risk effectively. For instance, for the standard mode, starting from inf TTC, and risk of 0, and then allows the other vehicle to turn. Yet its approach remains aggressive: by $4\mathrm{s}, TT$C falls to just 0.4s and risk spikes to 0.7, even though no collision occurs. Such behavior is risky and unsuitable for AV deployment. EthicAR w/o PER agent displays a similar pattern. The SACLAG agent shows improved restraint: though ego speed remains relatively high at $4\mathrm{s}, TT$C and risk increase to about 0.3, still yielding more courteously. By contrast, EthicAR maintains the lowest risk level throughout the scenario, embodying responsible and socially considerate driving.
Figure 8: Real‑world driving scenario in which the ego vehicle (green) and an opposing left‑turning vehicle (red) both prepare to turn across each other, with other traffic shown in gray. (a) Per‑timestep behavior of EthicAR and baseline agents, showing real‑time TTC with the cyclist and the maximum other risk of the ego vehicle. (b) TTC versus maximum risk at the five selected instants, plotted for direct quantitative comparison.